A certain radioactive material is known to decay at a rate proportional to the amount



present. If the initially there is 50 milligrams of the material present and after two



hours it is observed that the material has lost 10% of its original mass. Find



a) An expression for the mass of the material remaining at any time t



b) The mass of the material after 4 hours



c) The time rate at which the material has decayed to one half of its initial mass

1
Expert's answer
2022-02-01T16:46:47-0500

Let "A(t)=A=" the amount of radioactive material.

Given "\dfrac{dA}{dt}=-kA," where "k" is the proportionality constant.


"\dfrac{dA}{A}=-kdt"

Integrate


"\int\dfrac{dA}{A}=-\int kdt"

Let "A(0)=A_0=" the initialamount of radioactive material.

"A=A_0e^{-kt}"

Given "A_0=50\ mg, A(2)=0.9A_0"

Substitute


"0.9A_0=A_0e^{-k(2)}"

"-2k=\ln(0.9)"

"k=-\dfrac{1}{2}\ln (0.9) \ h^{-1}"


a)


"A(t)=50e^{\frac{1}{2}\ln (0.9)t}"

"A(t)=50(0.9)^{t\/2}"

b)


"A(4)=50(0.9)^{4\/2}"

"A(4)=40.5\ mg"

c)


"A(t_1)=\dfrac{1}{2}A_0"

"\dfrac{1}{2}A_0=A_0(0.9)^{t_1/2}"

"(0.9)^{-t_1\/2}=2"

"t_1=-\dfrac{2\ln2}{\ln 0.9}"

"t_1\approx13.1576\ hour"